### v1

# Scalable Dynamic Origin-Destination Demand Estimation Enhanced by High-Resolution Satellite Imagery Data

Abstract
This study presents a novel integrated framework for dynamic origin-destination demand estimation (DODE) in multi-class mesoscopic network models, leveraging high-resolution satellite imagery together with conventional traffic data from local sensors. Unlike sparse local detectors, satellite imagery offers consistent, city-wide road and traffic information of both parking and moving vehicles, overcoming data availability limitations. To extract information from imagery data, we design a computer vision pipeline for class-specific vehicle detection and map matching, generating link-level traffic density observations by vehicle class. Building upon this information, we formulate a computational graph-based DODE model that calibrates dynamic network states by jointly matching observed traffic counts and travel times from local sensors with density measurements derived from satellite imagery. To assess the accuracy and scalability of the proposed framework, we conduct a series of numerical experiments using both synthetic and real-world data. The results of out-of-sample tests demonstrate that supplementing traditional data with satellite-derived density significantly improves estimation performance, especially for links without local sensors. Real-world experiments also confirm the framework’s capability to handle large-scale networks, supporting its potential for practical deployment in cities of varying sizes. Sensitivity analysis further evaluates the impact of data quality related to satellite imagery data.
keywords:
Dynamic origin-destination demand estimation, remote sensing, high-resolution satellite imagery, computational graph, multi-source data
journal: Elsevier
1 Introduction
The widespread availability of spatio-temporal data has created new opportunities for advancing computational tools to model network flows, individual traveler behavior, and travel demand in dynamic transportation networks. Recent developments in sensing technologies and artificial intelligence are revolutionizing traditional models, making them more data-driven, scalable, and effective for complex, large-scale networks. Dynamic Origin-destination Demand Estimation (DODE) is a foundational prerequisite for dynamic network models to accurately reproduce the status quo spatio-temporal network conditions, supporting traffic assignment [Pi et al., 2019] and control strategies [Ye et al., 2019, Liu, Ma and Qian, 2023, Ke et al., 2025]. DODE studies can be broadly categorized into model-based methods, which embed physics-informed traffic assignment models, and model-free methods, which formulate the problem using data-driven techniques without traffic assignment constraints. Recent studies have merged the advantages of both approaches, reformulating DODE within learning-based frameworks and demonstrating promise in large-scale applications [Wu et al., 2018, Liu and Yin, 2025]. Despite extensive research over the past few decades, DODE remains challenging because it is inherently a high data-demanding problem due to its high-dimensional and under-deterministic nature. Accurate and reliable DODE, particularly in large-scale networks, requires the integration of diverse and comprehensive data sources.
Existing DODE studies reply on either single or multiple data sources from fixed traffic detectors (e.g., traffic count), mobile sensors (e.g., probe vehicle traces, speeds) and activity-based data (e.g., surveys, mobile device traces). However, one major limitation shared by these data sources is that they capture only a subset of travelers within the system due to the sparsity of sensor geographical coverage, low penetration rate of mobile sensors, low sample rate of survey and privacy issues of data collection. Prior studies demonstrated that the performance of DODE is highly sensitive to sensor placement [Zhu et al., 2016, 2017] while optimal sensor deployment requires long-term planning for new installations and routine maintenance for deployed sensors, both of which can be costly and time-consuming [Zhou and List, 2010, Salari et al., 2021, Li et al., 2023, Hu and Fan, 2024].
In recent years, satellite remote sensing data have emerged promising new data sources for city-wide transportation models, overcoming weakness of localized and mobile ground-based sensors [Sheehan et al., 2023, Gagliardi et al., 2023]. It has demonstrated effectiveness in monitoring both transportation infrastructure conditions and network-wide traffic dynamics. These data often have larger geographical coverage, and provide detailed, consistent information of all individual vehicles detected and road features in city-wide networks. Advances in computer vision and machine learning have significantly enhanced the information extraction process from these satellite images, including vehicle detection [Leitloff et al., 2010, Chen et al., 2014, Cao et al., 2016] and road segmentation [Dai et al., 2020, Mei et al., 2021]. Fusing this emerging data source, which is unconstrained by spatial locations, into DODE has great potential to make it more scalable, accurate and economic.
This study investigates the potential of incorporating high-resolution imagery data collected by flexible sky-view sensing techniques such as satellite into a computation graph-based DODE framework to enhance model scalability and performance in replicating real-world network-wise traffic conditions. The contributions are summarized as follows:
  • 1.
    We develop a novel dynamic OD demand estimation framework fusing multi-modal data including tabular data from localized detectors and imagery data from satellite. The satellite imagery data which is not restricted by the location, complementing the limitations of local sensors and having potentials for scalable applications.
  • 2.
    We develop a computer vision pipeline to extract information from high-resolution satellite imagery data, including detecting both moving vehicles and on-street parking vehicles by class, matching vehicles to links and computing link-level traffic densities which is a new observation in DODE.
  • 3.
    We design a general mesoscopic dynamic network loading (DNL) model compatible with traffic density estimation, and develop a DODE framework that integrates multi-source traffic condition data collected from local sensors and satellite images. The framework is formulated on computational graphs and solved efficiently on large-scale networks. Additionally, we address the data availability and inconsistency issues of multi-source data fusion.
  • 4.
    We use both synthetic and real-world data on toy and large-scale networks to evaluate model performance improvement when incorporating information from satellite imagery data. We perform sensitivity analysis to examine model robustness with respect to data quality, including computer vision sensing errors and imagery collection frequency.
The organization of the remaining paper is as follows: Section 2 summarizes related studies in DODE and usage of satellite imagery in transportation research. Section 3 introduces the proposed DODE framework integrating multi-source data including satellite imagery information and the general DNL model for density estimation. Section 4 conducts numerical experiments using both real-world multi-source data and synthetic data. Sensitivity analysis is also performed to test model’s robustness to data quality of density information extracted from satellite imagery. Section 5 summarizes findings and provides future research directions.
2 Related Work
2.1 Dynamic OD Demand Estimation
From a modeling perspective, DODE studies can be broadly categorized into model-based, model-free and overlapping approaches. Model-based approaches typically use a bi-level optimization structure, where upper level minimizes observed and estimated traffic conditions, and the lower level adopts a Dynamic Traffic Assignment (DTA) model to capture travel behavior and network traffic dynamics. In contrast, model-free methods do not use physical models and rely on data-driven approaches using various machine learning architectures. Overlapping approaches leverage advantages of both model-based and model-free approaches.
Regardless of the modeling approach, DODE problems are highly data-demanding. The traffic information data used in DODE are collected using either fixed detectors or mobile detectors. Fixed detectors has permanently fixed locations and provide continuous localized traffic conditions at specific locations. Some examples are traffic count data collected by loop detectors, automated vehicle identification (AVI) data collected by communication with the vehicle tags and barrier devices, and license plate recognition (LPR) data capture by fixed point monitoring cameras. Mobile detectors refer to portable device attached to a subset of travelers in the system and collect successive localized traffic conditions based on location changes of these travelers. Some examples are link speed data collected by probe vehicles on the link, complete vehicle traces of probe vehicles and human’s activity trajectories collected by mobile phone location tracking. Table 1 summarizes representative studies utilizing various data sources and modeling approaches.
Table 1: Representative studies
representative literature data type model type
Dixon and Rilett [2002] automated vehicle identification &\&& vehicle traces model-free
Zhou and Mahmassani [2006] automated vehicle identification data model-based
Frederix et al. [2010] traffic flow, link density model-based
Toledo and Kolechkina [2012] traffic count model-based
Wang et al. [2013] cell phone location data model-based
Lu et al. [2013] traffic count model-based
Lu et al. [2015] traffic count model-based
Carrese et al. [2017] traffic count, vehicle trajectories model-based
Ou et al. [2019] traffic count model-free
Mo et al. [2020] license plate recognition data model-based
Krishnakumari et al. [2020] traffic count model-free
Ma et al. [2020] traffic count &\&& speed model-based
Tang et al. [2021] automated vehicle identification data model-free
Cao et al. [2021] automated vehicle identification &\&& vehicle traces model-free
Ros-Roca et al. [2022] traffic count, link travel time model-free
Kumarage et al. [2023] traffic count model-based
Huo et al. [2023] traffic count model-based
Sun and Fan [2024] traffic count model-based
Lu et al. [2024] location based social network data model-based
Yumin Cao and Kang [2024] vehicle trajectories model-based
However, both fixed and mobile detectors have limitations. Fixed detector data relies heavily on sensor locations and often sparsely distributed in large networks, making it challenging to infer full network demand [Hu and Fan, 2024]. Additionally, maintaining installed sensors and planning new locations can be costly especially in large-scale networks, limiting scalability and sustainability. Mobile detectors, while more flexible, depends on penetration rates and a subset of travelers may not well represent the entire population [He et al., 2024]. Furthermore, both sensing categories suffer from inconsistent data quality at different times, different locations and different data sources, leading to contradictory conditions collected for the same link. To overcome these limitations, some studies have fused multiple complementary data sources, usually from different detectors, to improve DODE performance. Frederix et al. [2010] combined additional link-level density information into DODE, outperforming traditional estimation approaches using only traffic flow, which is not effective in discerning between congested and uncongested traffic states. Carrese et al. [2017] incorporate probe vehicle traces data with link traffic flow measurements to distinguish congestion states and enhance DODE reliability.
Recent studies have reformulated OD demand estimation and related problems within machine learning framework, proposing computational graph-based approaches that efficiently fuse multiple data sources. Wu et al. [2018] proposed a multi-layered computational graph representation to structurally model different levels of travel demand variables and individual travel behavior parameters. Ma et al. [2020] formulated the DODE problem on a computational graph utilizing multi-source multi-class traffic count and speed data. Lu et al. [2023] proposed an physics-informed framework for joint traffic state and queue profile estimation on layered computational graph by leveraging heterogeneous data sources. Kim et al. [2024] further developed a comprehensive analytical mathematical formulation to integrate transport demand and supply models in a computational graph framework. Guarda et al. [2024] developed a computational graph framework to estimate network flow and travel behavior using day-to-day system-level data and Guarda and Qian [2024] uses this data-driven framework to estimate traffic conditions in unobserved network locations. Liu and Qian [2024] estimated curbside usage demand patterns based on additional curb event monitoring data along with traffic count and speed information. Liu, Yin, Bai and Grimm [2023] and Liu and Yin [2025] embedded neural networks in implicit layers of an end-to-end learning framework to directly learns travel choice preferences, supply and demand components and the equilibrium state from multi-day link flow observations. Du et al. [2025] developed a fully differentiable end-to-end simulation-based optimization framework to calibrate passenger route choice using smart card and train loadings data. These studies demonstrated the potential of integrating diverse data sources, especially new data sources that can well address the concern of inconsistency information, within unified learning frameworks, leading to robust and precise demand estimation.
2.2 Satellite imagery in transportation research
Advances in computer vision and sensing has unlocked the potential of remote sensing data for enhancing city-wide infrastructure monitoring Gagliardi et al. [2023] and traffic condition monitoring [Sheehan et al., 2023]. High-resolution imagery data captures consistent, city-wide traffic conditions of all links at the same time, overcoming the limitations of ground-based sensors. Some pilot studies prove that the effectiveness of traffic monitoring using satellite imagery data [McCord et al., 2003, Larsen et al., 2009, 2013, Kaack et al., 2019]. Additionally, Unmanned Aerial Vehicles (UAV) can be another similar but more flexible approach to collect precise imagery or video-based traffic data in small sub-regions of city, due to their rapid speed, cost-effectiveness, and extensive field-of-view [Balamuralidhar et al., 2021, Butilă and Boboc, 2022, Bai and Feng, 2025].
Some pilot studies have explored the utilization of these imagery data for transportation research, including road network extraction [Dai et al., 2020, Mei et al., 2021], vehicle detection and classification [Leitloff et al., 2010, Chen et al., 2014, Cao et al., 2016]. Some studies have focused on traffic density estimation based on imagery or video collected from satellite [Kopsiaftis and Karantzalos, 2015, Sakai et al., 2019]. Reksten and Salberg [2021] proposed a deep learning-based framework to estimate annual average daily traffic (AADT) using very-high-resolution optical remote-sensing imagery with spatially limited ground-based measurements. Ganji et al. [2022] used Google aerial images together with sparse data from traffic counts to predict traffic volumes. Chen et al. [2021] used traffic density detection data to examine changes in spatio-temporal traffic patterns during the COVID 19 pandemic. He et al. [2021] integrated satellite imagery and GPS trajectories to generate high-resolution accident risk maps. Yu et al. [2024] integrated high-resolution satellite imagery with geo-spatial tabular and crash data to analyze factors contributing to freight truck-related crashes. Horsler et al. [2023] incorporated information from satellite imagery to predict regional road transport emissions, highlighting the potential in environmental impact assessments. Our initial work proves the potential of incorporating satellite imagery into mesoscopic network model calibration but the information extraction from imagery data is not vehicle class specific and the parking information is not considered, leading to limited accuracy in large-scale network applications [Liu et al., 2024].
To the best of our knowledge, none of the existing studies incorporated detailed traffic information of both moving and on-street parking vehicles by class extracted from satellite images directly into calibrating dynamic network models and performing DODE. Incorporating satellite imagery has the potential to fill gaps in areas where ground-based sensors are not available, providing a more complete picture of consistent network conditions to enhance DODE performance.
3 Model
This section introduces the DODE framework leveraging high-resolution satellite images as an additional data source. Figure 1 demonstrates overall framework consisting three parts: (1) multi-source data collection, including satellite imagery, road network and tabular data from localized ground-based sensors. (2) a computer vision pipeline to extract traffic information from satellite images and (3) DODE model incorporating all available data sources. In the following subsections, we first introduce the computer vision pipeline for density extraction. Next we formulate the DODE problem using traffic information from local ground-based detectors and satellite images with an introduction of a general dynamic network loading model. Furthermore, we present the framework on a computational graph to solve more efficiently using forward-backward algorithms. Lastly, we discuss data availability issues and corresponding aggregation methods.
Figure 1: Model Overview
3.1 Computer vision pipeline for traffic information extraction from satellite images
3.1.1 Vehicle detection
Vehicle detection in satellite imagery is a well-studied problem in computer vision, and various pre-trained models have been open-sourced [Nachmany and Alemohammad, 2019]. Liu et al. [2024] compared the performance of two model architectures provided in the Python deep learning library Raster Vision [Azavea/Element 84, Robert Cheetham, n.d.]. The first architecture is a Single Shot Multibox Detector (SSD) model [Liu et al., 2016], which uses ”VGG” as backbone [Simonyan and Zisserman, 2014]. The second architecture is a Faster R-CNN (FRCNN) model [Ren et al., 2017] that uses Resnet50 as backbone [He et al., 2016]. Liu et al. [2024] found that FRCNN gives the best results in terms of precision, recall and F1-score, and this study adopts the same Faster R-CNN architecture but extends its application to a multi-class classification setting, enabling the model to distinguish between cars and trucks.
The object detection model was trained using high-resolution satellite imagery from the xView dataset [Lam et al., 2018], which contains 744 annotated images at a 30-centimeter resolution. The dataset includes 60 object categories, 26 of which correspond to various types of vehicles. These 60 child categories are grouped into 7 parent classes. When labelers are unable to assign an object to a specific child category confidently, they fall back to the more general parent category [Lam et al., 2018]. Instead of classifying each object into its original xView category, the model was designed to perform a coarser multi-class classification task by categorizing detected objects as either car, truck, other vehicle, or non-vehicle. The car class includes the parent class Passenger Vehicle and the child class Small Car. The truck class includes the parent class Truck and nine truck-related child categories, such as Cargo Truck, Truck Tractor, and Haul Truck.
3.1.2 Map matching and density calculation
Raster Vision provides built-in functionality to extract the geospatial coordinates of the bounding boxes associated with cars and trucks in the satellite image. Using standard geospatial analysis tools, the centroids of these bounding boxes were computed and subsequently matched to the nearest links in the street network. To avoid assigning vehicles parking far from road segments or on segments not included in the network, only centroid points of vehicles falling within a predefined buffer around each road centerline were considered eligible for map matching.
After cars and trucks are matched to the links of the street network, traffic density for each link can be calculated as the ratio between the number of matched vehicles and the corresponding link length. This computation is performed separately for cars and trucks, as the DODE framework is capable of incorporating multi-class traffic density data. Because the link lengths are exogenous information, we use the total detected and matched vehicle number by class to represent link density without loss of generality.
3.2 Dynamic Origin-destination Demand Estimation (DODE)
This section formulates the DODE problem that integrates multi-source mobility data, including link-level traffic count, travel time and additional traffic density extracted from satellite images. We begin by describing the approach to modeling link-level densities for different vehicle classes in mesoscopic DNL model. Next, we formulate the DODE problem and present it on a computational graph and derive the forward and backward processes. Last, we discuss potential issues that may arise when deploying this framework in real-world scenarios with various data availabilities and the corresponding aggregation approaches.
3.2.1 Preliminary
Let $G=(N,A)$ represent a general road network with a node set $N$ and a link set $A$. A subset of links, denoted as $A_{p}\subset A$, represents links allowing for on-street parking. The assignment horizon is defined as $T_{d}=[1,2,...,T]$. $R\subset N$ and $S\subset N$ are sets of origin and destination nodes respectively. Between an origin-destination (OD) pair $rs$ where $r,s\in R,S$, $q^{rs}_{i,t}$ represents the total demand of vehicle class $i$ departing at time $t\in T_{d}$ from $r$ and heading to $s$. We consider two vehicle classes in this study: cars denoted by $\mathcal{C}$ and trucks denoted by $\mathcal{T}$. $P^{rs}_{i}$ is the set of paths available for vehicle class $i$ to choose between the OD pair $rs$, and $f^{rs}_{k,i,t}$ represents the path flow (number of trips) of vehicle class $i$ departing at time $t$ and choosing path $k\in P^{rs}_{i}$. The corresponding path travel cost is denoted as $c^{rs}_{i,k,t}$. $p^{rs}_{k,i,t}$ represents route choice proportion of choosing path $k$ for vehicle class $i$ departing from $r$ at time $t$. Therefore, path flow $f^{rs}_{k,i,t}$ can be obtained by multiplying the route choice proportion $p^{rs}_{k,i,t}$ with OD demand $q^{rs}_{i,t}$.
$f^{rs}_{k,i,t}=p^{rs}_{k,i,t}\cdot q^{rs}_{i,t},\ k\in P^{rs}_{i},\ rs\in RS,\ t\in T_{d}$ (1)
and in vectorized form
$\bm{f}_{i}=\bm{p}_{i}\bm{q}_{i},\ \forall i\in C$ (2)
The route choice portions $p^{rs}_{k,i,t}$ can be determined using various route choice function, such as dynamic user equilibrium [Pi et al., 2019, Liu and Qian, 2024] and stochastic user equilibrium based on Logit model [Guarda et al., 2024]. The inputs depend on the specific formulation and here we use path and link level travel costs, denoted by $c^{rs}_{k,i,t}$ and $h^{t}_{a,i}$, without loss of generality.
$p^{rs}_{k,i,t}=\Psi\left(\{c^{rs}_{k,i,t}\}_{rs,i,k,t},\{h^{t}_{a,i}\}_{i,a,t}\right)$ (3)
Vectorized form can be written as
$\bm{p}_{i}=\Psi(\{\bm{c}_{i}\}_{i},\{\bm{{h}}_{i}\}_{i}),\ \forall i\in C$ (4)
In this study, the agent-based mesoscopic dynamic network loading (DNL) model, MAC-POSTS [Pi et al., 2018, Ma et al., 2020], is used to simulate network-wide time-varying traffic conditions, including traffic count, travel time and density. In the DNL, cumulative curves are installed at the head and tail of each observed link to extract time-varying traffic conditions.
For links with traffic flow and/or travel time observations, we use two cumulative curves recording cumulative numbers of arriving and departing through vehicles respectively on link $a$, denoted by $\mathcal{A}^{m}_{a,i}$ and $\mathcal{D}^{m}_{a,i}$ respectively. Note that these vehicles do not stop (i.e., on-street parking) on this link and First-In-First-Out (FIFO) is enforced, hence they can be used to extract link travel times. For links with density observations, especially links in $A_{p}$ with on-street parking spaces, we install two additional cumulative curves at the head and tail of link $a$, denoted by $\mathcal{A}^{p}_{a,i}$ and $\mathcal{D}^{p}_{a,i}$ respectively, to record in/outflow for parking vehicles. Due to stochastic parking durations, FIFO is not enforced in this pair of cumulative curves and they are not used to extract travel time. Figure 2 demonstrates the setting of two pairs of cumulative curves at observed links with on-street parking spaces. To efficiently extract DAR matrices, these cumulative curves are tree-based [Ma et al., 2020, Liu and Qian, 2024].
(a) Cumulative curves for vehicles in through traffic
(b) Cumulative curves for on-street parking vehicles
Figure 2: Traffic condition extracted from cumulative curves
We incorporate curbside parking dynamics into the DNL model using the approach proposed in Liu and Qian [2024]. Truncated fundamental diagrams with respect to curb occupancies are used to govern link dynamics, which can be generalized to any link dynamics model, such as link queue or link transmission models. For instance, for link queue model, we can use link-level parking queues (instead of curb cells in CTM) to store parking vehicles, with the same curb parking dynamics to track curb availability and departure times.
Using these two pairs of cumulative curves, traffic count of class $i$ during time interval $t$ is modeled as the total arrival vehicle number of both parking and through vehicles during time interval $t$, shown as
$x^{t}_{\text{arr},a,i}=x^{t,m}_{\text{arr},a,i}+x^{t,p}_{\text{arr},a,i}=\mathcal{A}^{m}_{a,i}(\bar{t})-\mathcal{A}^{m}_{a,i}(\underline{t})+\mathcal{A}^{p}_{a,i}(\bar{t})-\mathcal{A}^{p}_{a,i}(\underline{t})$ (5)
Link travel time at time $t$ is modeled as the traversing time for vehicles arriving at time $t$ on link $a$, extracted from cumulative curves of through vehicles, which reads
$h^{t}_{a,i}=\mathcal{D}^{m,-1}_{a,i}(\mathcal{A}^{m}_{a,i}(t))-t=\bar{\Lambda}\left(x^{t}_{\text{arr},a,i}\right)$ (6)
where $\bar{\Lambda}$ is a general link dynamic function using link inflow as input because both arrival and departure cumulative curves can be seen as a function of link inflow, and link flow dynamics governs the relations between two curves.
3.2.2 Modeling traffic density
Traffic density is a measure of the number of vehicles per unit length of road. Because road length are exogenous, we use the total number of vehicles remaining on link $a$ at timestamp $\bar{t}$, denoted by $R_{a,i}(\bar{t})$ to represent the link density, which can be modeled as summation of remaining through and parking vehicle numbers at time $t$, each of them is the difference between cumulative arrival vehicle number and cumulative departure vehicle number before timestamp $t$. As our DNL model is discrete in time, $R_{a,i}(\bar{t})$ (the upper bound) is used to represent the density for time interval $t$.
$R_{a,i}(\bar{t})=\mathcal{A}^{m}_{a,i}(\bar{t})+\mathcal{A}^{p}_{a,i}(\bar{t})-\mathcal{D}^{m}_{a,i}(\bar{t})-\mathcal{D}^{p}_{a,i}(\bar{t})$ (7)
and the link density at time $\bar{t}$ is
$k^{{t}}_{a,i}=\frac{R_{a,i}(\bar{t})}{l_{a}}$ (8)
To obtain more stable and smooth density estimation, we average the densities across several neighboring time intervals around time ${t}$, denoted by $[{t}-\delta,{t}+\delta]$, the average density can be computed by
$k^{{t}}_{a,i}\cdot l_{a}=\frac{1}{2\delta+1}\sum_{t^{\prime}={t}-\delta}^{{t}+\delta}k^{t^{\prime}}_{a,i}\cdot l_{a}$ (9)
To make simulation-based DODE problem tractable, we linearize the DNL process using Dynamic Assignment Ratio (DAR) matrix [Ma et al., 2020], which mapping time-varying path flows to time-varying link flows. Because both arrival and departure link flows are used in estimating remaining vehicle numbers, and both parking and non-parking vehicles are accounted in link inflows, four separate DAR matrices are used to record corresponding mapping relationships. The traffic flow, represented as link inflow in our DNL, during time $t_{2}$ has the following relationship with path flows
$x^{t_{2}}_{\text{arr},a,i}=\sum_{r,s\in R,S}\sum_{k\in P^{rs}_{i}}\sum_{t_{1}\in T_{d}}\rho_{\text{arr},k,a,i}^{rs,m}(t_{1},t_{2})\cdot f^{rs}_{k,i,t_{1}}+\sum_{r,s\in R,S}\sum_{k\in P^{rs}_{i}}\sum_{t_{1}\in T_{d}}\rho_{\text{arr},k,a,i}^{rs,p}(t_{1},t_{2})\cdot f^{rs}_{k,i,t_{1}}$ (10)
where $\rho_{\text{arr},k,a,i}^{rs,n/p}(t_{1},t_{2})$ represents the time-dependent incidence matrix, which is the proportion of the path flow $f^{rs}_{k,i,t_{1}}$ contributing to the link arrival flow $x^{t_{2},n/p}_{\text{arr},a,i}$. Here we use $t_{1}$ and $t_{2}$ to differentiate time index for path and link flows respectively. The corresponding vectorized form can be written as
$\bm{x}_{i}=\bm{\rho}^{m}_{a,i}\bm{f}_{i}+\bm{\rho}^{p}_{a,i}\bm{f}_{i}$ (11)
Similarly, the departure flow of link $a$ at time $t_{2}$ for class $i$ can be modeled using departure DAR matrix as
$x^{t_{2}}_{\text{dep},a,i}=\sum_{r,s\in R,S}\sum_{k\in P^{rs}_{i}}\sum_{t_{1}\in T_{d}}\rho_{\text{dep},k,a,i}^{rs,m}(t_{1},t_{2})\cdot f^{rs}_{k,i,t_{1}}+\sum_{r,s\in R,S}\sum_{k\in P^{rs}_{i}}\sum_{t_{1}\in T_{d}}\rho_{\text{dep},k,a,i}^{rs,p}(t_{1},t_{2})\cdot f^{rs}_{k,i,t_{1}}$ (12)
Then the cumulative arrival and departure flows on link $a$ at time ${t}$ for vehicle class $i$ can be computed as the summation of all arrival or departure flows before time ${t}$, that are arrival or departure flows of intervals from $1$ to $t$,
$A_{a,i}(\bar{t})=\sum_{t_{2}=1}^{t}x^{t_{2}}_{\text{arr},a,i}\quad\text{and}\quad D_{a,i}(\bar{t})=\sum_{t_{2}=1}^{t}x^{t_{2}}_{\text{dep},a,i}$ (13)
Combining Equation 10,12 and 13, link density and path flow have the following linear relationship
$k^{t}_{a,i}\cdot l_{a}$ $=R_{a,i}(\bar{t})=A_{a,i}(\bar{t})-D_{a,i}(\bar{t})$ (14)
$=\sum_{t_{2}=1}^{t}\sum_{r,s\in R,S}\sum_{k\in P^{rs}_{i}}\sum_{t_{1}\in T_{d}}\rho_{\text{dep},k,a,i}^{rs,m}(t_{1},t_{2})\cdot f^{rs}_{k,i,t_{1}}+\sum_{t_{2}=1}^{t}\sum_{r,s\in R,S}\sum_{k\in P^{rs}_{i}}\sum_{t_{1}\in T_{d}}\rho_{\text{dep},k,a,i}^{rs,p}(t_{1},t_{2})\cdot f^{rs}_{k,i,t_{1}}$
$-\sum_{t_{2}=1}^{t}\sum_{r,s\in R,S}\sum_{k\in P^{rs}_{i}}\sum_{t_{1}\in T_{d}}\rho_{\text{dep},k,a,i}^{rs,m}(t_{1},t_{2})\cdot f^{rs}_{k,i,t_{1}}-\sum_{t_{2}=1}^{t}\sum_{r,s\in R,S}\sum_{k\in P^{rs}_{i}}\sum_{t_{1}\in T_{d}}\rho_{\text{dep},k,a,i}^{rs,p}(t_{1},t_{2})\cdot f^{rs}_{k,i,t_{1}}$
The corresponding vectorized form is
$\bm{k}_{i}\odot\bm{l}=\bm{H}_{i}(\bm{\rho}^{m}_{a,i}\bm{f}_{i}+\bm{\rho}^{p}_{a,i}\bm{f}_{i})-\bm{H}_{i}(\bm{\rho}^{m}_{d,i}\bm{f}_{i}+\bm{\rho}^{p}_{d,i}\bm{f}_{i})$ (15)
where $\bm{H}_{i}$ is used to aggregate total arrival and departure flow from time interval $1$ to $t_{2}$ in Equation 14.
To summarize, the DNL model, denoted by $\Lambda$, uses time-varying path flows as inputs and gives out link level traffic conditions and DAR matrices mapping recording the mapping relations between path flows and link flows, shown in Equation (16).
$\left\{x^{t_{2}}_{\text{arr},a,i},h^{t_{2}}_{a,i},k^{t_{2}}_{a,i},\rho_{\text{arr},k,a,i}^{rs,m}(t_{1},t_{2}),\rho_{\text{arr},k,a,i}^{rs,p}(t_{1},t_{2}),\rho_{\text{dep},k,a,i}^{rs,m}(t_{1},t_{2}),\rho_{\text{dep},k,a,i}^{rs,p}(t_{1},t_{2})\right\}_{rs,i,k,a,t_{1},t_{2}}$ (16)
$=\Lambda\left(\{f^{rs}_{k,i,t}\}_{rs,k,i,t}\right)$
In vectorized form, the DNL function can be written as
$\{\bm{x}_{i},\bm{{h}}_{i},\bm{k}_{i},\bm{\rho}^{m}_{a,i},\bm{\rho}^{p}_{a,i},\bm{\rho}^{m}_{d,i},\bm{\rho}^{p}_{d,i}\}_{i}=\Lambda(\{\bm{f}_{i}\}_{i}),\ \forall i\in C$ (17)
3.2.3 DODE formulation
The DODE can be formulated as a mathematical programming problem shown in Equation 18. The decision variable is OD demand $\{\bm{q}_{i}\}_{i}$. The objective function is minimizing the differences between modeled and observed traffic conditions including link traffic count, link travel time and link density. The constraints include route choice, DNL, path-demand relationship and non-negativity constraint of demand.
$\min_{\{\bm{q}_{i}\}_{i}}$ $\mathscr{L}=\ w_{1}{\left(\left\|\bm{x}^{o}-\sum_{i}\bm{L}_{i}\bm{{x}}_{i}\right\|^{2}_{2}\right)}+w_{2}{\left(\left\|\bm{h}^{o}-\sum_{i}\bm{M}_{i}\bm{{h}}_{i}\right\|^{2}_{2}\right)}+w_{3}{\left(\left\|\bm{k}^{o}\odot\bm{l}-\sum_{i}\bm{I}_{i}\bm{{k}}_{i}\odot\bm{l}\right\|^{2}_{2}\right)}$ (18)
[…]
$\bm{q}_{i}\geq 0,\ \forall i\in C$
where $\bm{x}^{o}$, $\bm{h}^{o}$ and $\bm{k}^{o}\odot\bm{l}$ are observed count, travel time and remaining vehicles (density multiplied by link length). $\bm{L}_{i}$, $\bm{M}_{i}$ and $\bm{I}_{i}$ are aggregation matrices used to map simulated conditions to available observations, depending on data availability which will be discussed in section 3.4.
3.3 Computational graph-based solution algorithm
The formulation can be presented on a computational graph and solved by forward-backward algorithms efficiently. In the forward pass shown in Figure 3, a dynamic traffic assignment problem is solved using behavior function $\Psi$ and DNL model $\Lambda$, given fixed time-varying multi-class demand $\{\bm{q}_{i}\}_{i}$ (initialized demand or estimated demand from previous epoch). Route choices $\{\bm{p}_{i}\}_{i}$, link-level traffic states and DAR matrices $\bm{\rho}^{m}_{a,i},\bm{\rho}^{p}_{a,i},\bm{\rho}^{m}_{d,i},\bm{\rho}^{p}_{d,i}$ are obtained from Equations (17), (4) and (2). To improve computation efficiency in DNL on large-scale networks, we can use hybrid route choice models which is embedded in the DNL model to retrieve link traffic measurements and DAR matrices in one DNL run. In the hybrid route choice model, a fraction of travelers follow pre-trip route choice model and the other travelers adaptively explore time-varying shortest paths following en-route route choice model [Qian and Zhang, 2013, Liu and Qian, 2024].
In the backward pass, shown in Figure 3, the gradient of the loss function with respect to demand is derived using the Chain Rule in Equation 19. The gradient of objective function with respect to decision variable $\bm{q}_{i}$ is shown in Equation 20. Any gradient-based solution algorithm can be used to solve the DODE based on this gradient on the computational graph.
$\frac{\partial\mathscr{L}}{\partial\mathscr{L}_{1}}$ $=w_{1}$ (19)
$\frac{\partial\mathscr{L}_{1}}{\partial\bm{x}_{i}}$ $=-2\bm{L}_{i}^{T}\left(\bm{x}^{o}-\sum_{i\in C}\bm{L}_{i}\bm{x}_{i}\right)$
$\frac{\partial\mathscr{L}_{1}}{\partial\bm{f}_{i}}$ $=({\bm{\rho}^{m}_{a,i}}^{\intercal}+{\bm{\rho}^{p}_{a,i}}^{\intercal})\frac{\partial\mathscr{L}_{1}}{\partial\bm{x}_{i}}$
$\frac{\partial\mathscr{L}_{1}}{\partial\bm{q}_{i}}$ $=\bm{p}^{T}_{i}\frac{\partial\mathscr{L}_{1}}{\partial\bm{f}_{i}}$
$\frac{\partial\mathscr{L}}{\partial\mathscr{L}_{2}}$ $=w_{2}$
$\frac{\partial\mathscr{L}_{2}}{\partial\bm{h}_{i}}$ $=-2\bm{M}_{i}^{T}\left(\bm{h}^{o}-\sum_{i\in C}\bm{M}_{i}\bm{h}_{i}\right)$
$\frac{\partial\mathscr{L}_{2}}{\partial\bm{f}_{i}}$ $=\frac{\partial\bm{h}_{i}}{\partial\bm{f}_{i}}\frac{\partial\mathscr{L}_{2}}{\partial\bm{h}_{i}}$
$\frac{\partial\mathscr{L}_{2}}{\partial\bm{q}_{i}}$ $=\bm{p}^{T}_{i}\frac{\partial\mathscr{L}_{2}}{\partial\bm{f}_{i}}$
$\frac{\partial\mathscr{L}}{\partial\mathscr{L}_{3}}$ $=w_{3}$
$\frac{\partial\mathscr{L}_{3}}{\partial\bm{r}_{i}}$ $=-2\bm{I}^{T}_{i}\left(\bm{k}^{o}\odot\bm{l}-\sum_{i}\bm{I}_{i}\bm{k}_{i}\odot\bm{l}\right)$
$\frac{\partial\mathscr{L}_{3}}{\partial\bm{f}_{i}}$ $=\left({\bm{\rho}^{m}_{a,i}}^{\intercal}+{\bm{\rho}^{p}_{a,i}}^{\intercal}-{\bm{\rho}^{m}_{d,i}}^{\intercal}-{\bm{\rho}^{p}_{d,i}}^{\intercal}\right)\bm{H}_{i}^{\intercal}\frac{\partial\mathscr{L}_{3}}{\partial\bm{r}_{i}}$
$\frac{\partial\mathscr{L}_{3}}{\partial\bm{q}_{i}}$ $=\bm{p}^{T}_{i}\frac{\partial\mathscr{L}_{3}}{\partial\bm{f}_{i}}$
$\frac{\partial\mathscr{L}}{\partial\bm{q}_{i}}=$ $-2w_{1}\bm{p}_{i}^{\intercal}({\bm{\rho}^{m}_{a,i}}^{\intercal}+{\bm{\rho}^{p}_{a,i}}^{\intercal})\bm{L}_{i}^{\intercal}\left(\bm{x}^{o}-\sum_{i}\bm{L}_{i}\bm{x}_{i}\right)-2w_{2}\bm{p}^{\intercal}_{i}\frac{\partial{\bm{{h}}_{i}}}{\partial\bm{f}_{i}}\bm{M}_{i}^{\intercal}\left(\bm{h}^{o}-\sum_{i}\bm{M}_{i}\bm{{h}}_{i}\right)$ (20)
$-2w_{3}\bm{p}^{\intercal}_{i}\cdot\left({\bm{\rho}^{m}_{a,i}}^{\intercal}+{\bm{\rho}^{p}_{a,i}}^{\intercal}-{\bm{\rho}^{m}_{d,i}}^{\intercal}-{\bm{\rho}^{p}_{d,i}}^{\intercal}\right)\bm{H}_{i}^{\intercal}\bm{I}_{i}^{\intercal}\left(\bm{k}^{o}\odot\bm{l}-\sum_{i}\bm{I}_{i}\bm{k}_{i}\odot\bm{l}\right)$
3.4 Data availability and aggregation
This section addresses issues related to data availability and aggregation methods. The DNL model simulates network-wide traffic conditions in various resolutions of interest (in this study, we estimate traffic conditions of each link every 15 minutes), however, observed data may not be collected at the same frequency. Moreover, data quality and pre-processing error cannot provide precise spatio-temporal features. To address this, we use aggregation matrices to mapping modeled conditions to observations and aggregation can occur in both spatial and temporal domains.. Temporal aggregation refers to aggregating traffic conditions into larger time intervals. For example, summing up four 15min traffic count to match hourly traffic flow observation [Ma et al., 2020, Liu and Qian, 2024].
Spatial aggregation primarily arises from limitations in satellite image quality and computer vision method’s performance. We discuss three potential scenarios which require spatial aggregation. The first scenario is about lane-level vehicle detection. when CV pipeline does not have precise road segmentation feature, it cannot conduct lane-level vehicle detection, failing to differentiate between parking and through vehicles. In such cases, observations provide link density for all vehicles on the link and we aggregate parking and through vehicles together to model density, which is already described in Equation (7) and (8). In this study, we integrate the aggregation into density estimation in the DNL model. Another general way is estimating separate parking density and through traffic density in the DNL and using matrix $\bm{I}_{i}$ to aggregate them. Another reason we aggregating them in DNL process is that one through vehicle captured by satellite image only represent the status at the timestamp of the image and we are not sure if this vehicle will perform on-street parking after this timestamp or just left a parking spot before this timestamp. Aggregating through and parking vehicles into total density can overcome this information gap. The second scenario is about vehicle moving directions. Similar to scenario 1, if CV pipeline cannot perform road segmentation and detect the vehicle orientations, detected vehicles cannot be correctly matched to directed road segments, potentially leading to flow inconsistency on successive links. To address this, we aggregate two (even more) moving directions together to match regional total density. Another scenario is about vehicle class detection. When the CV pipeline fails to detect vehicle classes with high precision, we can simply aggregate density by vehicle classes.
(a) Forward process
(b) Backward process
Figure 3: Computational graph of DODE
4 Numerical Experiments
4.1 Toy network with synthetic data
We begin with a synthetic toy network to evaluate the effectiveness of incorporating density observations into DODE. The network, shown in Figure 4 is modified from Zhang and Qian [2020] which consists of 18 links, including 12 road segments and 6 OD connectors. Detailed link configurations an be found in Table 5. Time-varying traffic conditions, including traffic counts, travel times and densities, are generated based on synthetic demand and recorded at 15-minute intervals to serve as ground truth. As the description in Section 3, we aggregate moving and parking vehicles together to calculate the total remaining vehicles around the segment to represent link densities. To simulate measurement noise, traffic count and travel time observations are corrupted by multiplying by random error terms $\epsilon_{10\%}$ drawn from uniform distribution $\text{Unif}(0.9,1.1)$. The densities are also corrupted by random multiplicative errors of three levels with corresponding uniform distribution: (1) ±$\pm 10\%$ with $\epsilon_{10\%}\sim\text{Unif}(0.9,1.1)$ (2) ±$\pm 20\%$ with $\epsilon_{20\%}\sim\text{Unif}(0.8,1.2)$ and (3) ±$\pm 50\%$ with $\epsilon_{50\%}\sim\text{Unif}(0.5,1.5)$.
To simulate the sparsity of real-world observations, 6 out of 12 links are randomly selected to form the observed link set, denoted by $A_{obs}$. The remaining 6 links are out-of-sample (OOS) links, denoted by link set $A_{oos}$. To prove incorporating density snapshots of the whole network can enhance overall DODE performance, we conduct two separate DODE tests under two data scenarios and compare estimation accuracy on both $A_{obs}$ and $A_{oos}$. In scenario 1, we use only traffic counts and travel times from $A_{obs}$, represented as $\{x_{a}^{t},h_{a}^{t}\}_{t\in T_{d},a\in A_{obs}}$ to mimic the DODE process using localized sensor data. In scenario 2, the data is augmented with 10 density snapshots of all 12 links, which are taken at 15-minute intervals (same as traffic count and travel time data), denoted by $\{k_{a}^{t}\}_{t\in T_{d},a\in A}$. We use coefficient of determination, denoted by $R^{2}$, to measure estimation accuracy of traffic conditions throughout this paper. The main difference of this test between Liu et al. [2024] is incorporating parking information in the link density observations.
Figure 5, 6, 7 and 8 show the count and travel time estimation for unobserved and observed links under different scenario settings. It can be seen that augmenting data with traffic density information for all links can improve the estimation accuracy of both unobserved and observed links, therefore improving the overall performance. However, the improvement relies on the sensing error level when extracting density information from imagery data. The improvement for OOS links are overall better than observed links since density is the only data source for their estimation. Even though the test is performed on a toy network, the DODE is still under-deterministic because only half of links has traffic count and speed observations and the density information cannot differentiate parking and moving vehicles, and these information cannot fully reproduce the truth OD matrix. We calculate the mean squared errors for demand estimation under different scenario settings (shown in Table 2), demonstrating incorporating density can lead to estimated demand which is more closer to ground truth and the improvement of car demand estimation is better than truck due to the relatively higher car volumes on network.
Figure 4: Small network
(a) Scenario 1
(b) Scenario 2 (±$\pm 10\%$ error in densities)
(c) Scenario 2 (±$\pm 20\%$ error in densities)
(d) Scenario 2 (±$\pm 50\%$ error in densities)
Figure 5: Comparison of count estimation for out-of-sample links
(a) Scenario 1
(b) Scenario 2 (±$\pm 10\%$ error in densities)
(c) Scenario 2 (±$\pm 20\%$ error in densities)
(d) Scenario 2 (±$\pm 50\%$ error in densities)
Figure 6: Comparison of travel time estimation for out-of-sample links
(a) Scenario 1
(b) Scenario 2 (+/- 10% error in densities)
(c) Scenario 2 (+/- 20% error in densities)
(d) Scenario 2 (+/- 50% error in densities)
Figure 7: Comparison of count estimation for observed links
(a) Scenario 1
(b) Scenario 2 (+/- 10% error in densities)
(c) Scenario 2 (+/- 20% error in densities)
(d) Scenario 2 (+/- 50% error in densities)
Figure 8: Comparison of travel time estimation for observed links
(a) Scenario 1
(b) Scenario 2 (+/- 10% error in densities)
(c) Scenario 2 (+/- 20% error in densities)
(d) Scenario 2 (+/- 50% error in densities)
Figure 9: Comparison of density estimation for all links
Vehicle class Scenario 1 Scenario 2 (±$\pm 10\%$) Scenario 2 (±$\pm 20\%$) Scenario 2 (±$\pm 50\%$)
car 34.4 26.5 28.3 28.3
truck 2.9 2.4 2.4 2.7
Table 2: Mean absolute error (MAE) of multi-class demand under different scenarios
4.2 Pittsburgh Downtown network with real-world data
4.2.1 Experiment Setting
The numerical experiment uses real-world and synthetic data on Pittsburgh Downtown network to evaluate the proposed framework. The network and sensing scales are shown in Figure 10 and detailed configuration is presented in Table 6. Two distinct real-world data sources are used: traffic count collected by detectors and manual counting provided by the Southwestern Pennsylvania Commission, and two high-resolution satellite images provided by XX. The proposed computer vision pipeline is applied to extract link densities from the satellite images, distinguishing between vehicle classes but not parking and through vehicles. Therefore the resultant link density observations represent total vehicles of both parking and through traffic.
Figure 10: Pittsburgh network
4.2.2 Traffic density observations from satellite imagery
Two 30-cm resolution satellite images captured by a Pleiades Neo satellite from Airbus were used as the input for the computer vision pipeline. The images were collected on Wednesday, May 11th, 2022 at 12:18 PM and Monday, March 6th, 2023 at 12:17 PM and in Eastern Time. As shown in the first column of images in Figure 12, each image has a width of 5km and a height of approximately 4 kilometers, spanning around 20 square kilometers of surface and covering most of the downtown area in Pittsburgh, PA. The original images included four channels, but only the RGB channels were extracted to run the computer vision pipeline.
To improve object detection results with Airbus imagery, the object detector is also trained with 30-cm resolution satellite imagery from the xView dataset [Lam et al., 2018]. Using satellite imagery with the same resolution for training and inference has been shown to improve object detection results [Van Etten, 2018]. To ensure adequate representation of the positive classes, only the top 10 percent of images with the highest number of annotated vehicles were selected. This subset was then divided into training and validation sets, consisting of 46 images (80%) and 13 images (20%), respectively.
Table 3 presents the precision, recall, and F1-score on the training set, disaggregated by vehicle type. The recall and precision for cars are comparable to those reported by Liu et al. [2024], which used a binary classifier for car detection. The lower precision observed for trucks relative to cars aligns with the trend in mean average precision reported by Lam et al. [2018] for truck-related classes in the xView dataset. The last five columns of the table report the observed counts, predicted counts, true positives, false positives, and false negatives for each vehicle type, where the predicted count equals the sum of true and false positives, and the observed count equals the sum of true positives and false negatives.
Vehicle type Metric Precision Recall F1 score Observed count Predicted count True positives False positives False negatives Cars $0.62$ $0.66$ $0.64$ $15,562$ $16,496$ $10,256$ $6,240$ $5,306$ Trucks $0.33$ $0.31$ $0.32$ $1,177$ $1,127$ $369$ $758$ $808$ Cars and trucks $0.60$ $0.63$ $0.62$ $16,739$ $17,623$ $10,625$ $6,998$ $6,114$
Table 3: Performance metrics for object detection in validation set
Figure 11 shows vehicle detection, classification and matching results for image patches with a height and width of 100 meters. To analyze results across different types of streets, image patches were collected in areas containing local roads and highways. The yellow bounding boxes in the second column show the vehicle detection output. The third column shows the classification results, with purple and blue boxes representing cars and trucks, respectively. In the fourth column, the red points are the vehicles matched to the street network, while the yellow points represent vehicles that were detected but not matched to the street networks. To exclude vehicles far from the road segments in the network, only vehicles near the road centerline are matched to the street network. Therefore, as shown in the image of the local road in the bottom right corner, vehicles parked near buildings or houses are not matched to the street network.
Figure 11: Vehicle detection, classification, and matching results using satellite image collected on May 11, 2022. ……………………….
Figure 12 shows the traffic density at each link of the street network disaggregated by trucks and cars, with the left column showing the satellite images collected for each year. Links with the darkest red color represent traffic densities higher than 40 vehicles per lane-kilometer and are associated with high congestion levels. The traffic density patterns for cars and trucks are consistent across images, which is expected given that both images were collected during the same hour of the day and on weekdays. In addition, the estimated traffic density is higher for cars than for trucks, which is in line with the share of cars and trucks observed in real-world transportation networks.
Figure 12: Link-level traffic density derived from satellite imagery, disaggregated by year (2022, 2023) and vehicle class (truck, car)………
After computer vision processing, two class-specific density snapshots are obtained within the same time interval (12:15-12:30 PM) on different dates. Figure 23 illustrates the differences between the two data samples. To mitigate the inconsistency, a threshold-based approach is adopted, where links with a difference of at most 5 are selected as consistent links, and their densities are used in the experiments. As a result, 900 out of 1044 links are considered to be consistent, shown as red dots in Figure 23.
To address the inconsistency introduced by multi-source data (e.g., different data providers and different sensing errors in data collection), we propose a two-stage calibration process using two satellite images separately. The first stage calibration is performed using one density snapshot and traffic count data, and we select a subset of links based on the following criteria: for links with non-zero density observations, we selected the link with estimated density no more than three times the observed value for both vehicle classes; for links with zero density observations, we select link with estimated densities no more than 10. As a result, 457 out of 900 link are selected for the second stage. In the second stage, we use the second densities of the selected link subset to perform a separate calibration together with the same traffic count data. This two-stage calibration process helps mitigate the consistency between data sources and leads to more reliable results.
4.2.3 Results and discussion
To evaluate performance of adding additional density data in DODE, we design two DODE scenarios similar to toy network, but this time we only have information of observed links and the ground truth of unobserved links remains hidden. Scenario 1 uses only localized traffic count and speed data of observed links and scenario 2 uses both local sensors and network-wide density snapshot. As DODE is a widely acknowledged underestimate problem with unknown true OD demand and traffic states of unobserved links, we propose the following conditions to be satisfied, justifying the claim that incorporating satellite imagery can enhance DODE performance: (1) adding density into DODE cannot do harm to other data sources, meaning that the accuracy of other traffic condition estimation cannot be largely reduced. (2) the accuracy of density estimation should be improved for observed links, i.e., larger R-squared values for link with density observations. (3) the estimated OD demand pattern in scenario 2 is different from the one in scenario 1.
Figure 13 shows the convergence curves for both DODE tests of two scenarios. Both curves converge after around 70 epochs and the total reduction is similar, indicating that adding additional density loss term in the objective function of DODE will not do harm to the convergence performance. We further plot the estimation results of link-level conditions. Figure 14 shows the accuracy of estimated conditions in DODE without density. It can be seen that link flow estimations match observations well with R-squared values of above 0.9 but using the same network OD flow pattern cannot reproduce the densities. However, when we incorporate density observations in the DODE, the results (Figure 15) show that the estimated flows still match observed traffic counts (with a slight reduction in accuracy) and more importantly, the estimated density can match with observations with much higher accuracy. This can be explained by the under-deterministic nature of DODE problem. Utilizing partial observations (i.e., local sensor data) cannot guarantee the uniqueness of OD matrix. The different demand patterns might lead to the same local traffic conditions, resulting in highly biased estimation, which is difficult to recognize. Adding more data source (i.e., density) is adding more constraints in the problem, leading to smaller feasible region and the estimation results can be closer to the true values. We plot the estimated OD patterns of these two DODE runs and observe some discrepancies for some OD pairs in Figure 16 and 17. The estimation results with density observation satisfy the proposed conditions and we can say incorporating additional density observations can enhance DODE performance.
We further investigate how adding density observations will impact the accuracy of the same localized traffic flow estimation due to the ”trade-off” of fitting multi-source data. We focus on ten overlapping links with both traffic count and density observations and plot the estimation results in two scenarios. Because the satellite image only provide one density snapshot which is in the time interval of 12:15-12:30 PM, we compare the traffic flow estimation in this interval, shown in Figures 18 and 19. It shows that the link flow estimations do not have large discrepancies between two scenarios while truck flow estimations performs worse than cars especially for observed flows that are smaller than 10. This is due to the stochasticity in DNL model and relatively low detection accuracy for trucks in CV pipeline. The unknown inconsistencies between data sources also contribute to this difference.
(a) DODE without density
(b) DODE with density
Figure 13: Convergence curve
(a) link flow estimation
(b) density estimation
Figure 14: DODE without density data
(a) link flow estimation
(b) density estimation
Figure 15: DODE with density data
Figure 16: OD demand results from two scenario (car)
Figure 17: OD demand results from two scenario (truck)
Figure 18: Car flow estimation of overlapping links
Figure 19: Truck flow estimation of overlapping links
4.3 Pittsburgh Downtown network with synthetic data
Given the limited coverage of real-world traffic data in Section 4.2.1, this experiment uses synthetic data to conduct a sensitivity analysis and evaluate the robustness of the proposed DODE framework incorporating density observations. We use the estimated multi-class OD demand from Scenario 2 in Section 4.2.1(which incorporates both traffic count and density snapshots) as the ”ground truth” OD demand. This demand is fed into the DNL model to generate time-varying traffic conditions across all links, serving as the synthetic ground truth observations.
Following the data generation process described in Section 4.1, we randomly sample 400 links to form the observation link set $A_{obs}$, and introduce stochastic errors to traffic count, speed and density data to simulate sensing errors. The remaining links form the out-of-sample link set $A_{oos}$ for which only density observations with stochastic errors are available. These corrupted observations, including traffic count, speed of observed links and density on all links, are used as input for DODE.
We evaluate the accuracy of estimated traffic conditions separately for observed and unobserved links. This section focuses on analyzing the sensitivity of the DODE framework to varying data conditions, specially the level of sensing error and frequency of imagery-based density snapshots. To reduce the bias introduced by spatial distribution of local sensors, each DODE test is repeated with five different random samples of $A_{obs}$ under each experimental setting.
We first assess the sensitivity of DODE to sensing errors in density observations. As discussed in Section 3.4, density observations may require aggregation when serving as the inputs for DODE and aggregating moving and through vehicles together can introduce sensing errors. Additionally, inaccuracies in class-specific vehicle detection further contribute to noises in sensing results.
To evaluate the impact of such errors, this section considers two levels of sensing errors: (1) ±$\pm 10\%$ with perturbation $\epsilon_{10\%}\sim\text{Unif}(0.9,1.1)$ (2) ±$\pm 20\%$ with perturbation $\epsilon_{20\%}\sim\text{Unif}(0.8,1.2)$. For each error level, 5 DODE runs are conducted using independently sampled observed link sets ($A_{\text{obs}}$) to mitigate the impact related to spatial sampling of local sensors. Consistent with prior sections, $R^{2}$ is used to evaluate model performance. To assess model generalization, we report separate $R^{2}$s for observed and unobserved links to assess the model generalization.
Figure 20 shows the a clear trend that increasing sensor error level from ±$\pm 10\%$ to ±$\pm 20\%$ leads to a noticeable decline in $R^{2}$ values, indicating estimation accuracy is reduced. The flow estimates for unobserved links are relatively more robust, which maintains similar performance under both error level settings, but for density estimation, $R^{2}$ deteriorates with higher error level because sensing errors can reduce the consistency in network-wide density observations. These results demonstrate that the sensing errors in density extraction from imagery data do not have large impact on the local traffic estimation at unobserved locations but will impact overall density estimation itself.
(a) observed car flow estimation
(b) observed truck flow estimation
(c) unobserved car flow estimation
(d) unobserved truck flow estimation
(e) observed car density estimation
(f) observed truck density estimation
Figure 20: Sensitivity analysis for density sensing error
Next sensitivity analysis is evaluating the impact of capture frequency of satellite imagery. We fix the same level of sensing errors in density observations as ±$\pm 10\%$ and change density snapshot frequencies in two different levels: (1) density snapshots per 15 min and (2) density snapshots per 30 min. Similarly, 5 DODE runs are conducted in each data setting and separate $R^{2}$s for observed and unobserved links are reported to assess model performance.
(a) observed car flow estimation
(b) observed truck flow estimation
(c) unobserved car flow estimation
(d) unobserved truck flow estimation
(e) observed car density estimation
(f) observed truck density estimation
Figure 21: Sensitivity analysis for density snapshot frequency
Figure 21 shows that across all metrics, estimation accuracy, as measured by R-squared, tends to decline as the density snapshot frequency decreases from 15 to 30 minutes This indicates that denser temporal coverage contributes to more accurate calibration.The impact is more pronounced for truck flow estimation than for car flow, as truck volumes are generally lower and thus more sensitive to observational constraints. For density estimations, the 30-min snapshot frequency results in noticeably poorer performance compared to the 15-minute case, likely due to increased uncertainty in the aggregated density observations. Higher snapshot frequency captures more detailed spatiotemporal traffic conditions, which improves estimation accuracy. Similar trends can be observed for car and truck flow estimation on unobserved (out-of-sample) links, suggesting that more frequent density snapshots enhance estimation accuracy for links lacking local sensor data. Nevertheless, estimation accuracy for unobserved links remains lower than that for observed links, reflecting the under-deterministic nature of the DODE problem.
5 Conclusion
This study proposes a scalable computational graph-based framework to estimate dynamic origin-destination demand by integrating multi-modal data including emerging high-resolution satellite imagery and other conventional mobility data collected from ground-based sensors. Satellite imagery data provides city-wide class-specific traffic density information for both on-road and parking vehicles, which is extracted using computer vision techniques, overcoming the limitations of local ground-based traffic sensor data. A series of numerical experiments using both synthetic and real-world data on networks of various sizes demonstrate that incorporating satellite imagery-derived density can improve estimation performance of the whole network, especially for segments without local sensor coverage. We also conduct sensitivity analysis in terms of sensing error and snapshot frequency in imagery-derived density observations to evaluate the robustness of our model to various data availability conditions.
Future work can be summarized as the following directions: from data perspective, more city-wide satellite imagery data can be collected and used for testing model robustness in real-world applications. For modeling parking behavior, satellite imagery only captures the conditions of curbside spaces and outdoor parking lots. Fusing other parking information such as indoor parking condition and transactions of parking meters will enhance more comprehensive parking choices. From modeling perspective, the computer vision pipeline can be enhanced with advanced lane-level road segmentation features which can differentiate parking and moving vehicles and facilitate more accurate map matching. The computer vision can also be improved for accurate vehicle classification. The proposed model can also be extended to use video-based data which provides traffic information of higher resolution.
Acknowledgment
This research is sponsored by Fujitsu Research of America.
Appendix
Notation Description
$G$ road network
$N$ node set
$A$ link set
$A_{p}$ link set allowed for on-street parking
$T_{d}$ assignment time horizon
$R$ origin node set
$S$ destination node set
$q^{rs}_{i,t}$ demand of vehicle class $i$ departing at time $t$ from $r$ and heading to $s$
$\mathcal{C}$ car class
$\mathcal{T}$ truck class
$P^{rs}_{i}$ path set for vehicle class $i$ between $rs$
$f^{rs}_{k,i,t}$ flow on path $k$ for vehicle class $i$ departing at time $t$ between OD pair $rs$
$c^{rs}_{k,i,t}$ travel cost of path $k$ for vehicle class $i$ departing at time $t$ between OD pair $rs$
$p^{rs}_{k,i,t}$ route choice proportion of choosing path $k$ for vehicle class $i$ departing at time $t$ between OD pair $rs$
$h^{t}_{a,i}$ travel time of link $a$ for vehicle class $i$ at time $t$
$\Psi$ route choice function
$\mathcal{A}^{m}_{a,i}$ arrival curve of link $a$ for moving vehicles of class $i$
$\mathcal{A}^{p}_{a,i}$ arrival curve of link $a$ for parking vehicles of class $i$
$\mathcal{D}^{m}_{a,i}$ departure curve of link $a$ for moving vehicles of class $i$
$\mathcal{D}^{p}_{a,i}$ departure curve of link $a$ for parking vehicles of class $i$
$x^{t}_{\text{arr},a,i}$ total arrival flow on link $a$ for vehicle class $i$ at time $t$
$x^{t,m}_{\text{arr},a,i}$ arrival flow on link $a$ for moving vehicle of class $i$ at time $t$
$x^{t,p}_{\text{arr},a,i}$ arrival flow on link $a$ for parking vehicle of class $i$ at time $t$
$\bar{\Lambda}$ link dynamic function
$R_{a,i}(t)$ remaining vehicle number of class $i$ on link $a$ at time $t$
$l_{a}$ length of link $a$
$k^{t}_{a,i}$ density of link $a$ for vehicle class $i$ at time $t$
$\rho^{rs,m}_{\text{arr},k,a,i}(t_{1},t_{2})$ the proportion of $k$-th path flow departing at time $t_{1}$ that arrives on link $a$ at time $t_{2}$ recorded in cumulative curves for moving vehicles
$\rho^{rs,p}_{\text{arr},k,a,i}(t_{1},t_{2})$ the proportion of $k$-th path flow departing at time $t_{1}$ that arrives on link $a$ at time $t_{2}$ recorded in cumulative curves for parking vehicles
$\rho^{rs,m}_{\text{dep},k,a,i}(t_{1},t_{2})$ the proportion of $k$-th path flow departing at time $t_{1}$ that departs from link $a$ at time $t_{2}$ recorded in cumulative curves for moving vehicles
$\rho^{rs,p}_{\text{dep},k,a,i}(t_{1},t_{2})$ the proportion of $k$-th path flow departing at time $t_{1}$ that departs from link $a$ at time $t_{2}$ recorded in cumulative curves for parking vehicles
$A_{a,i}(\bar{t})$ total cumulative arrival flow before timestamp $\bar{t}$ of link $a$ for class $i$
$D_{a,i}(\bar{t})$ total cumulative departure flow before timestamp $\bar{t}$ of link $a$ for class $i$
$\Lambda$ dynamic network loading function
$\bm{f}_{i}$ vectorized path flow of class $i$
$\bm{k}_{i}$ vectorized link density of class $i$
$\bm{q}_{i}$ vectorized demand of class $i$
$\bm{l}$ vectorized link length of class $i$
$\bm{\rho}^{m}_{a,i}$ vectorized DAR matrix for arrival moving vehicles of class $i$
$\bm{\rho}^{p}_{a,i}$ vectorized DAR matrix for arrival parking vehicles of class $i$
$\bm{\rho}^{m}_{d,i}$ vectorized DAR matrix for departure moving vehicles of class $i$
$\bm{\rho}^{p}_{d,i}$ vectorized DAR matrix for departure parking vehicles of class $i$
$\bm{x}_{i}$ vectorized link flow of class $i$
$\bm{h}_{i}$ vectorized link travel time of class $i$
$\bm{p}_{i}$ vectorized route choice proportion of class $i$
$\bm{c}_{i}$ vectorized travel cost of class $i$
$\bm{L}_{i}$ aggregation matrix for link flow observation of class $i$
$\bm{M}_{i}$ aggregation matrix for travel time observation of class $i$
$\bm{I}_{i}$ aggregation matrix for link density observation of class $i$
$\bm{H}_{i}$ aggregation matrix for summing total cumulative arrival and departure flow
$\bm{x}^{o}$ observed link flow
$\bm{h}^{o}$ observed link travel time
$\bm{k}^{o}$ observed link density
Table 4: Notations
ID length car free-flow speed truck free-flow speed car capacity truck capacity car jam density truck jam density
1 1 50 40 6000 3600 540 240
2 0.5 50 40 2000 1200 90 40
3 1 50 40 6000 3600 540 240
4 0.5 50 40 2000 1200 180 80
5 1 50 40 6000 3600 540 240
6 0.5 50 40 2000 1200 90 40
7 1 50 40 4000 2400 360 160
8 0.5 50 40 2000 1200 90 40
9 1 50 40 6000 3600 540 240
16 2 30 20 2000 1200 360 160
17 0.5 30 20 2000 1200 90 40
18 2 30 20 2000 1200 360 160
Table 5: Summary of link properties of small network
Name Value
Study period 11:00-13:00
Simulation time resolution 5s
Time interval for condition estimation 15min
Number of time intervals 8
Number of links 3548
Number of nodes 1515
Number of origins/destinations 126
Number of OD pairs 15876
Table 6: Pittsburgh network setting
Figure 22: Coverage of two satellite images from Airbus
Figure 23: Density comparison in two images